OpenStudy (anonymous):

what is the slope of the graph 2y-5x=14

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

What is your understanding of this Q? ^

OpenStudy (skullpatrol):

Have learned about the "slope-intercept" form of a line?

OpenStudy (anonymous):

you have to let x=0

OpenStudy (skullpatrol):

You don't need to do that for what this question is asking :-)

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

tha

OpenStudy (anonymous):

than not really idk

OpenStudy (anonymous):

need

OpenStudy (anonymous):

help

OpenStudy (skullpatrol):

the slope-intercept form is y = mx + b where m is the slope

OpenStudy (anonymous):

the slope is 14

OpenStudy (anonymous):

?

OpenStudy (skullpatrol):

Put 2y - 5x = 14 into the form y = mx + b then m will be the slope that the question is asking for.

OpenStudy (anonymous):

do i divide 5 into 2

OpenStudy (anonymous):

im new at this soo sorry

OpenStudy (skullpatrol):

2y - 5x = 14 2y = 5x + 14 y = ( 5/2)x + 14/2 y = m x + b What is m?

OpenStudy (anonymous):

2.5

OpenStudy (skullpatrol):

Correct! That is the slope of the graph of 2y - 5x = 14.

OpenStudy (anonymous):

how do i put it in it said i was wrong

OpenStudy (skullpatrol):

Check that you copied the right question.

OpenStudy (anonymous):

yeah same question same how i put it in

OpenStudy (skullpatrol):

Use 5/2

OpenStudy (anonymous):

that work thank you for your help

OpenStudy (skullpatrol):

Thanks for asking :-)